8 pens in a bag, 3 blue, 5 red. 2 taken out at random, without replacement. Probability they are the same colour?

Either both blue or both red. 'OR' rule - add probabilities.Blue - first blue and second blue - 'AND' rule - multiply probabilities. 3/8 x 2/7 (as not replaced) = 6/56 = 3/28 (multiply top and bottom)Red - 5/8 x 4/7 = 20/56 = 5/143/28 + 5/14 = 3/28 + 10/28 = 13/28

AS

Related Maths GCSE answers

All answers ▸

Solve the simultaneous equations. 2x + y = 18 . x - y = 6


How do I find f'(x) for f(x)=4x^3+x^2+5x+8?


Text books are stored on two shelves. Each shelf is 0.72m long. Each textbook is 30 millimetres wide. Could 50 textbooks be stored on these shelves? (3 marks)


The line L passes through the points (-2,3) and (6,9). How do I find the equation of the line that is parallel to L and passes through the point (5,-1)?